Due to changes necessitated by the COVID 19 outbreak, the Mavericks have made changes to their operations. They need to know how many boxes they will need to rent to make their desired weekly profits from luxury boxes. You are a Maryville Sports Business Management graduate brought in to improve the profitability of the stadium. Assume that the new costs are as follows:  variable costs will be $1,000 per luxury box rented and fixed costs will be $15,000. Assume the team wants to make $47,000 per week from rentals. Assume they charge $7,000 to rent each box (and partial rentals are not allowed.) How many boxes must they rent per week to earn their desired profit?
Answer #1

Variable costs per box = $1,000

Fixed costs = $15,000

Target profit = $47,000

Rent per box = $7,000

Contribution margin per unit = Rent per box - Variable costs per box

= 7,000 - 1,000

= $6,000

Units to be sold to get a target profit = (Fixed cost + Target profit)/Contribution margin per unit

= (15,000 + 47,000)/6,000

= 62,000/6,000

= 11 (rounded to next whole number)

They must rent per week to earn their desired profit = 11 boxes

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.